# Perspective on a challenge: predicting the photochemistry of cyclobutanone

### 14 Apr 2026 v1

@@ 1 Introduction
Developing methods to simulate the dynamics of electronically excited molecules beyond the Born–Oppenheimer approximation, coined nonadiabatic molecular dynamics, has been a very active field of research over the past decades with numerous applications, e.g., in spectroscopy,Liu et al. (2025) photochemistry,Levine and Martínez (2007); Schuurman and Stolow (2018) or atmospheric chemistry.Curchod and Orr-Ewing (2024) Nonadiabatic molecular dynamics relies not only on strategies that can capture the subtle coupling between electronic and nuclear motion in regions of strong nonadiabaticity, but also on the possibility to approximate some of the nuclear quantum effects unraveled by the photoexcitation process. Recent applications of nonadiabatic molecular dynamics to realistic simulations of molecules also highlighted the key importance of the underlying electronic-structure methods,Janoš and Slavíček (2023); Papineau, Jacquemin, and Vacher (2024); Jíra, Janoš, and Slavíček (2024) which provide electronic energies, nuclear forces, nonadiabatic coupling matrix elements, and other electronic quantities required for the nuclear propagation. Photophysical processes, where the nuclear configuration space visited by the photoexcited molecule does not include bond breakage or formation, can often be well described by simple, single-reference quantum-chemical methods. Conversely, photochemical processes, which span a much larger region of the nuclear configuration space, typically require a multireference description of the electronic wavefunctions. The methods picked for the nonadiabatic dynamics and electronic structure are, however, not the only important ingredients for a successful simulation of excited-state processes: the description of the photoexcitation process and the calculation of experimental observables of interest are key components for a direct comparison with experiments. Hence, performing nonadiabatic molecular dynamics requires control over a vast set of parameters for different highly entangled components: photoexcitation, electronic structure, nonadiabatic dynamics, and calculation of observables.Cigrang et al. (2025); Prlj et al. (2026)
@@ 2 Analysis of the contributions
Electronic-structure methods used for the challenge comprise single-reference methods such as MP2 (second-order Møller–Plesset theory), ADC(2) (algebraic diagrammatic construction to second order) or EOM-CCSD (equation-of-motion coupled cluster singles doubles), and density-based approaches such as DFT (density functional theory) and LR-TDDFT/TDA (linear-response time-dependent density functional theory with or without the Tamm-Dancoff approximation). Multiconfigurational methods used include SA-CASSCF (state-averaged complete active space self-consistent field), MCSCF (multiconfigurational self-consistent field), and FOMO-CASCI (floating occupation molecular orbital complete active space configuration interaction). The multireference strategies employed are MRCIS (multireference configuration interaction singles) and XMS-CASPT2 (extended multi-state complete active space second order perturbation theory). The semiempirical approach MRCI/ODM3 (multireference configuration interaction with the semiempirical orthogonalization and dispersion corrections) was also tested.
@@ 2.2 Electronic structure
| Label | Ref | ES method¶ | ES family | Basis set† | ES code |
| Paper 1⊗ | 20 | SR | def2-TZVPD | Turbomole | |
| Paper 2 | 21 | MR | aDZ | OpenMolcas | |
| Paper 3⊗ | 22 | SR | aDZ | eT | |
| Paper 4 | 23 | MR | def2-SVPD | Bagel | |
| Paper 5 | 24 | MC | aDZ | OpenMolcas | |
| Paper 6⊗ | 25 | _0 | MC | aDZ | Molpro |
| Paper 7⊗ | 26 | MC | 6-311++G** | Molpro | |
| Paper 8 | 27 | MR | 6-31++G* | OpenMolcas | |
| Paper 9 | 28 | MC | aDZ | Molpro | |
| Paper 10⊗ | 29 | MC | H: DZ; C: aDZ; O: d-aDZ | Columbus | |
| Paper 11⊗ | 30 | SR | aDZ | Turbomole | |
| Paper 12 | 14 | _0 | MR | aDZ | Bagel |
| Paper 13 | 31 | MC | 6-311+G* | TeraChem | |
| Paper 14 | 32 | MR | H: DZ; C/O: aDZ | Columbus | |
| Paper 15⊗ | 33 | MC | aDZ | OpenMolcas |
@@
Other contributions made use of multiconfigurational methods for their dynamics. In the following, we mention the active-space size (and state averaging procedure) – the reader can consult the corresponding references for details about the orbitals included. SA-CASSCF was used in several contributions with various combinations of active spaces and state averaging procedures (with or without triplet states included): SA3-CASSCF(12/12)/aug-cc-pVDZ (Paper 5 and Paper 9), SA3-CASSCF(12/11)/aug-cc-pVDZ followed by SS-CASSCF(12/11)/aug-cc-pVDZ for ground-state dynamics when encountering instabilities (Paper 6), SA3-CASSCF(8/10)/6-311++G∗∗ (Paper 7, for DD-vMCG with a database, see Sec. 2.3), and SA6-CASSCF(8/11)/aug-cc-pVDZ (Paper 15).
Multireference methods were represented predominantly by XMS-CASPT2: XMS(4S,4T)-CASPT2(8/8)/aug-cc-pVDZ (Paper 2), XMS3-CASPT2(8/10)/def2-SVPD (Paper 4), CASPT2 corrected SA(3S,3T)-CASSCF(10/8)/6-311++G∗∗ for linear vibronic coupling model used with ML-MCTDH (Paper 7), XMS(3S,3T)-CASPT2(8/8)/6-31++G∗ (Paper 8), XMS3-CASPT2(8/8)/aug-cc-pVDZ combined with MP2/aug-cc-pVDZ for ground-state dynamics when dissociation occurred (Paper 12).
@@
Isolating critical geometries (minima, transition states, minimum-energy conical intersections) on potential energy surfaces revealed that cyclobutanone can evolve adiabatically in S2, i.e., switching its electronic character from Rydberg to valence, with an associated bond elongation (carbon-carbon bond next to the carbonyl group) leading eventually to ring opening and nonadiabatic transitions from S2 to S1 first, and then to S0. The adiabatic evolution away from the region of Rydberg character involves a transition state, whose barrier height with respect to the S minimum (Fig. 4) is critical for the timescales depicted by nonadiabatic dynamics simulations (see Sec. 2.3). Perhaps more importantly, moving away from the Franck–Condon region revealed the limitations of single-reference methods and LR-TDDFT. Different contributions used interpolated or intrinsic-coordinate pathways between critical geometries to compare electronic energies along these paths as obtained by single-reference (and LR-TDDFT) and multiconfigurational/multireference methods. Such a benchmark beyond the Franck–Condon region highlights the multireference character and, as such, the potential issues that EOM-CCSD or LR-TDDFT can have beyond the description of the initially formed Rydberg state (Paper 12). But these pathways also make explicit the dramatic impact played by dynamic correlation on the S2 barrier controlling the adiabatic evolution from Rydberg to valence character (Paper 5, Paper 6, Paper 8, Paper 12). SA-CASSCF calculations predict an adiabatic evolution in S2 that is essentially barrierless, whereas including dynamic correlation (for example with XMS-CASPT2) leads to the appearance of a small yet non-negligible barrier (certain contributions isolated a transition state).
@@
Overall, the photochemistry of cyclobutanone initiated by a laser pulse at 200 nm offers a fantastic playground for testing electronic-structure strategies for various processes. Upon photoexcitation, the molecule undergoes a photophysical process by relaxing in S2 with Rydberg (3s) character, leading to the appearance of vibronic progressions in the photoabsorption cross-section of cyclobutanone. This process is accurately captured by single-reference and multireference methods as long as an adequate basis set is employed. However, cyclobutanone can escape this S minimum via a transition state, leading to an adiabatic change of electronic character and triggering its rich photochemistry. This second process is far more demanding for electronic-structure methods. Not only should the method be able to describe bond-breaking processes (via a multiconfigurational or multireference description of the electronic wavefunctions) but it should capture the barrier height properly to adequately describe the timescale for the adiabatic process to take place. The contributions to this challenge allow us to highlight the critical role of the electronic-structure method (in particular, the inclusion of both static and dynamic correlation to describe the adiabatic evolution in S2) and the basis set, but more calculations are required to fully characterize the barrier on this adiabatic path. While comparison of experimental and calculated lifetimes for specific features in the MeV-UED observables (see Sec. 2.4 below) may offer an indirect indication that specific approaches (in particular, XMS-CASPT2 with a large basis set) describe this barrier adequately, the uncertainty related to this barrier height in S2 can only be reduced by additional calculations at a higher level of theory, possibly including CC3 if the multireference character of this nuclear configuration is not too important. Discussions during the CECAM workshop pointed out that larger active spaces, in particular when combined with a perturbation-based description of dynamic correlation, may not provide a better estimate of the barrier height (see for example the SI of Ref. 37 for a discussion on active-space size with SA-CASSCF).
@@ 2.3 Nonadiabatic dynamics
Nonadiabatic molecular dynamics approaches combined with a multiconfigurational or multireference method (and not restricted by the definition of model Hamiltonian) all showed that, upon the adiabatic transformation of the S2 electronic state from Rydberg to valence character, cyclobutanone ring-opens and reaches an intersection with S1. Focusing on the population traces, it appears that SA-CASSCF-based nonadiabatic dynamics reach a higher S1 population than XMS-CASPT2-based simulations. Interpolated pathways comparing these regions indicate that the intersection seam between S2 and S1 is closer to S0 with XMS-CASPT2 than with SA-CASSCF, for the same active space and basis set (see for example Paper 12). A concern raised during the CECAM workshop was that higher-lying 3p Rydberg states may decrease in energy during the dynamics and along the adiabatic transformation in S2, potentially becoming relevant to the dynamics. In the Franck–Condon region, a large energy gap exists between the absorption band assigned to the S2(3s)S0 transition and those characteristic of higher singlet states. Consequently, most studies restricted their simulations to the three lowest singlet states (S0, S1, S2), with the exception of Papers 2, 3, and 15 (Paper 12 analyzed the potential importance of higher excited states in its Supplementary Material). Single-reference approaches such as LR-TDDFT facilitate the inclusion of a larger number of excited states. LR-TDDFT-based nonadiabatic dynamics in Paper 15 revealed a non-negligible S3 population, while Paper 3 also incorporated S3 and observed population transfer to this state, albeit to a much smaller extent. Including additional Rydberg states is particularly demanding for active-space-based methods and, as noted earlier, might be highly sensitive to the choice of basis set.
@@ 2.4 Calculation of observables
This comparative analysis allows us to stress some of the final findings of this exercise. First of all, most contributions employing a multiconfigurational or multireference method offer in general a good description of most features of the , yet with significant differences in the evolution timescales for the main positive and negative features of this observable. XMS-CASPT2-based nonadiabatic molecular dynamics simulations lead, overall, to a better agreement with experiment for the overall trends and lifetime of features appearing on the map (e.g., predicted signals in Fig. 7 for Papers 2, 8, 12). Interestingly, some inconsistency between multiconfigurational methods can be observed: predictions 7, 10 and 15, do not exhibit the same features as the rest of the predictions employing multiconfigurational methods (for Paper 15, instabilities in the active space were noted). This observation further highlights that multiconfigurational methods, namely SA-CASSCF in the present case, are not black box tools and specific settings of parameters, such as active space or number of states, strongly determine the outcome of the calculations. The predictions based on single-reference methods (predicted signals in Fig. 7 for Papers 1, 3, 11) mostly exhibit a nearly static signal emanating from dynamics near the S2(3s) excited-state minimum, lacking the dissociation features observed with predictions using a multireference/multiconfigurational method. Despite the apparent discrepancies between the predicted signals represented as colormaps, the reader should bear in mind that the quantitative analysis offered by most contributions (and depicted in Fig. 8 with lineouts) is already useful to understand the combination of processes occurring during the photodynamics of cyclobutanone.
@@
While the maps plotted in Fig. 7 are useful to depict the overall time-evolution of the UED signal, they do not permit a clear evaluation of the temporal evolution of their respective features. To unravel the time-dependence of key features in the signal, we provide lineouts at and in Fig. 8, using both absolute (Figs. 8A and C) and normalized (Figs. 8B and D) intensities. All lineouts are integrated in a range of to reduce experimental noise and mimic the standard analysis of an experimental signal. The normalized data (normalized with respect to the maximum value along each lineout) are provided here only to facilitate a comparison with the experimental data, but the comparison between absolute scale and normalized data highlights the danger of such a comparison (see end of this paragraph). The lineout at is connected to the ring opening of cyclobutanone, while that at is connected to the fragmentation of the molecule and formation of photoproducts. Fig. 8B reinforces the observation of the previous paragraph that predictions employing a multiconfigurational method predict too fast dynamics, whereas those making use of a multireference approach provide a closer match with the experimental time scale. This dramatic impact of the electronic-structure method on the outcome of the nonadiabatic molecular dynamics is clearly highlighted by Fig. 9A (and Fig. 9C), which represents the same lineouts as in Fig. 8 but now with each contribution colorcoded based on the family of electronic-structure method used (single-reference, multiconfigurational, multireference). Fig. 9A clearly highlights the trends (discussed earlier for the ring-opening process with a few exceptions): single-reference methods (yellow curves) lead to a slower adiabatic ring-opening process in S2 in comparison to nonadiabatic dynamics using multireference approaches (purple curves); multiconfigurational methods (cyan curves), on the contrary, induce a much faster ring-opening dynamics than that observed with multireference methods (except for the result of Paper 14). However, a correct timescale for a feature of the signal does not mean a correct prediction: Fig. 8A shows that the predictions can have significant variations in signal intensities, highlighting an important variation of the overall population of trajectories leading to a given signal. For example, the time trace of the normalized lineout at predicted by Paper 10 using MCSCF appears to have a reasonable timescale in Fig. 8B, but this feature is associated with a nearly zero intensity in the absolute signal (Fig. 8A). The initial increase of the experimental signal at (Fig. 8C and Fig. 8D) is due to the bond-elongation event during ring opening and dissociations. The following negative signal at later times is caused by the formation of photoproducts, which lack the signal associated with larger bond distances observed in cyclobutanone between carbons and oxygen. This feature is captured only by three predictions, namely Paper 6 (SA-CASSCF), 12 (XMS-CASPT2), and 14 (MRCIS), which simulated the ground-state dynamics of cyclobutanone and its photoproducts for a longer timescale. Finally, and keeping in mind the biased distribution towards mixed quantum/classical methods during this challenge, it is interesting to note that the trends observed in the lineouts do not seem to be correlated with the family of methods used for the nonadiabatic molecular dynamics (Fig. 9B and Fig. 9D). Additional simulations would be required to further support this lack of correlation.

### 16 Apr 2026 v2

@@ 1 Introduction
Developing methods to simulate the dynamics of electronically excited molecules beyond the Born–Oppenheimer approximation, coined nonadiabatic molecular dynamics, has been a very active field of research over the past decades with numerous applications, e.g., in spectroscopy,Liu et al. (2025) photochemistry,Levine and Martínez (2007); Schuurman and Stolow (2018) or atmospheric chemistry.Curchod and Orr-Ewing (2024) Nonadiabatic molecular dynamics relies not only on strategies that can capture the subtle coupling between electronic and nuclear motion in regions of strong nonadiabaticity, but also on the possibility to approximate some of the nuclear quantum effects unraveled by the photoexcitation process. Recent applications of nonadiabatic molecular dynamics to realistic simulations of molecules also highlighted the key importance of the underlying electronic-structure methods,Janoš and Slavíček (2023); Papineau, Jacquemin, and Vacher (2024); Jíra, Janoš, and Slavíček (2024) which provide electronic energies, nuclear forces, nonadiabatic coupling matrix elements, and other electronic quantities required for the nuclear propagation. Photophysical processes, where the nuclear configuration space visited by the photoexcited molecule does not include bond breakage or formation, can often be well described by simple, single-reference quantum-chemical methods. Conversely, photochemical processes, which span a much larger region of the nuclear configuration space, typically require electronic-structure methods including both static and dynamic correlation. The methods picked for the nonadiabatic dynamics and electronic structure are, however, not the only important ingredients for a successful simulation of excited-state processes: the description of the photoexcitation process and the calculation of experimental observables of interest are key components for a direct comparison with experiments. Hence, performing nonadiabatic molecular dynamics requires control over a vast set of parameters for different highly entangled components: photoexcitation, electronic structure, nonadiabatic dynamics, and calculation of observables.Cigrang et al. (2025); Prlj et al. (2026)
@@ 2 Analysis of the contributions
Electronic-structure methods used for the challenge comprise single-reference methods such as MP2 (second-order Møller–Plesset theory), ADC(2) (algebraic diagrammatic construction to second order) or EOM-CCSD (equation-of-motion coupled cluster singles doubles), and density-based approaches such as DFT (density functional theory) and LR-TDDFT/TDA (linear-response time-dependent density functional theory with or without the Tamm-Dancoff approximation). These single-reference methods only account for dynamic correlation. Methods that effectively only incorporate static correlation and used in this challenge include SA-CASSCF (state-averaged complete active space self-consistent field), MCSCF (multiconfigurational self-consistent field), and FOMO-CASCI (floating occupation molecular orbital complete active space configuration interaction). These methods will collectively be labeled as ’MCSCF’ in the following. Some contributions used electronic-structure methods incorporating both static and dynamic correlation, specifically MRCIS (multireference configuration interaction singles) and XMS-CASPT2 (extended multi-state complete active space second order perturbation theory). In the following, these methods including both static and dynamic correlation will be labeled with ’MRPT/MRCI’. The semiempirical approach MRCI/ODM3 (multireference configuration interaction with the semiempirical orthogonalization and dispersion corrections) was also tested.
@@ 2.2 Electronic structure
| Label | Ref | ES method¶ | Correlation | Basis set† | ES code |
| Paper 1⊗ | 20 | Dynamic | def2-TZVPD | Turbomole | |
| Paper 2 | 21 | Static + dynamic | aDZ | OpenMolcas | |
| Paper 3⊗ | 22 | Dynamic | aDZ | eT | |
| Paper 4 | 23 | Static + dynamic | def2-SVPD | Bagel | |
| Paper 5 | 24 | Static | aDZ | OpenMolcas | |
| Paper 6⊗ | 25 | _0 | Static | aDZ | Molpro |
| Paper 7⊗ | 26 | Static | 6-311++G** | Molpro | |
| Paper 8 | 27 | Static + dynamic | 6-31++G* | OpenMolcas | |
| Paper 9 | 28 | Static | aDZ | Molpro | |
| Paper 10⊗ | 29 | Static | H: DZ; C: aDZ; O: d-aDZ | Columbus | |
| Paper 11⊗ | 30 | Dynamic | aDZ | Turbomole | |
| Paper 12 | 14 | _0 | Static + dynamic | aDZ | Bagel |
| Paper 13 | 31 | Static | 6-311+G* | TeraChem | |
| Paper 14 | 32 | Static + dynamic | H: DZ; C/O: aDZ | Columbus | |
| Paper 15⊗ | 33 | Static | aDZ | OpenMolcas |
@@
Other contributions made use of MCSCF methods for their dynamics. In the following, we mention the active-space size (and state averaging procedure) – the reader can consult the corresponding references for details about the orbitals included. SA-CASSCF was used in several contributions with various combinations of active spaces and state averaging procedures (with or without triplet states included): SA3-CASSCF(12/12)/aug-cc-pVDZ (Paper 5 and Paper 9), SA3-CASSCF(12/11)/aug-cc-pVDZ followed by SS-CASSCF(12/11)/aug-cc-pVDZ for ground-state dynamics when encountering instabilities (Paper 6), SA3-CASSCF(8/10)/6-311++G∗∗ (Paper 7, for DD-vMCG with a database, see Sec. 2.3), and SA6-CASSCF(8/11)/aug-cc-pVDZ (Paper 15).
MRPT/MRCI methods were represented predominantly by XMS-CASPT2: XMS(4S,4T)-CASPT2(8/8)/aug-cc-pVDZ (Paper 2), XMS3-CASPT2(8/10)/def2-SVPD (Paper 4), CASPT2 corrected SA(3S,3T)-CASSCF(10/8)/6-311++G∗∗ for linear vibronic coupling model used with ML-MCTDH (Paper 7), XMS(3S,3T)-CASPT2(8/8)/6-31++G∗ (Paper 8), XMS3-CASPT2(8/8)/aug-cc-pVDZ combined with MP2/aug-cc-pVDZ for ground-state dynamics when dissociation occurred (Paper 12).
@@
Isolating critical geometries (minima, transition states, minimum-energy conical intersections) on potential energy surfaces revealed that cyclobutanone can evolve adiabatically in S2, i.e., switching its electronic character from Rydberg to valence, with an associated bond elongation (carbon-carbon bond next to the carbonyl group) leading eventually to ring opening and nonadiabatic transitions from S2 to S1 first, and then to S0. The adiabatic evolution away from the region of Rydberg character involves a transition state, whose barrier height with respect to the S minimum (Fig. 4) is critical for the timescales depicted by nonadiabatic dynamics simulations (see Sec. 2.3). Perhaps more importantly, moving away from the Franck–Condon region revealed the limitations of single-reference methods and LR-TDDFT. Different contributions used interpolated or intrinsic-coordinate pathways between critical geometries to compare electronic energies along these paths as obtained by single-reference (and LR-TDDFT) and MCSCF or MRPT/MRCI methods. Such a benchmark beyond the Franck–Condon region highlights the importance of static correlation and, as such, the potential issues that EOM-CCSD or LR-TDDFT can have beyond the description of the initially formed Rydberg state (Paper 12). But these pathways also make explicit the dramatic impact played by dynamic correlation on the S2 barrier controlling the adiabatic evolution from Rydberg to valence character (Paper 5, Paper 6, Paper 8, Paper 12). SA-CASSCF calculations predict an adiabatic evolution in S2 that is essentially barrierless, whereas including dynamic correlation (for example with XMS-CASPT2) leads to the appearance of a small yet non-negligible barrier (certain contributions isolated a transition state).
@@
Overall, the photochemistry of cyclobutanone initiated by a laser pulse at 200 nm offers a fantastic playground for testing electronic-structure strategies for various processes. Upon photoexcitation, the molecule undergoes a photophysical process by relaxing in S2 with Rydberg (3s) character, leading to the appearance of vibronic progressions in the photoabsorption cross-section of cyclobutanone. This process is accurately captured by single-reference and MRPT/MRCI methods that incorporate dynamic correlation, as long as an adequate basis set is employed. However, cyclobutanone can escape this S minimum via a transition state, leading to an adiabatic change of electronic character and triggering its rich photochemistry. This second process is far more demanding for electronic-structure methods. Not only should the method be able to describe bond-breaking processes (via the inclusion of static correlation) but it should capture the barrier height properly to adequately describe the timescale for the adiabatic process to take place. The contributions to this challenge allow us to highlight the critical role of the electronic-structure method (in particular, the inclusion of both static and dynamic correlation to describe the adiabatic evolution in S2) and the basis set, but more calculations are required to fully characterize the barrier on this adiabatic path. While comparison of experimental and calculated lifetimes for specific features in the MeV-UED observables (see Sec. 2.4 below) may offer an indirect indication that specific approaches (in particular, XMS-CASPT2 with a large basis set) describe this barrier adequately, the uncertainty related to this barrier height in S2 can only be reduced by additional calculations at a higher level of theory, possibly including CC3 if static correlation at this nuclear configuration is not too important. Discussions during the CECAM workshop pointed out that larger active spaces, in particular when combined with a perturbation-based description of dynamic correlation, may not provide a better estimate of the barrier height (see for example the SI of Ref. 37 for a discussion on active-space size with SA-CASSCF).
@@ 2.3 Nonadiabatic dynamics
Nonadiabatic molecular dynamics approaches combined with an MCSCF or MRPT/MRCI method (and not restricted by the definition of model Hamiltonian) all showed that, upon the adiabatic transformation of the S2 electronic state from Rydberg to valence character, cyclobutanone ring-opens and reaches an intersection with S1. Focusing on the population traces, it appears that SA-CASSCF-based nonadiabatic dynamics reach a higher S1 population than XMS-CASPT2-based simulations. Interpolated pathways comparing these regions indicate that the intersection seam between S2 and S1 is closer to S0 with XMS-CASPT2 than with SA-CASSCF, for the same active space and basis set (see for example Paper 12). A concern raised during the CECAM workshop was that higher-lying 3p Rydberg states may decrease in energy during the dynamics and along the adiabatic transformation in S2, potentially becoming relevant to the dynamics. In the Franck–Condon region, a large energy gap exists between the absorption band assigned to the S2(3s)S0 transition and those characteristic of higher singlet states. Consequently, most studies restricted their simulations to the three lowest singlet states (S0, S1, S2), with the exception of Papers 2, 3, and 15 (Paper 12 analyzed the potential importance of higher excited states in its Supplementary Material). Single-reference approaches such as LR-TDDFT facilitate the inclusion of a larger number of excited states. LR-TDDFT-based nonadiabatic dynamics in Paper 15 revealed a non-negligible S3 population, while Paper 3 also incorporated S3 and observed population transfer to this state, albeit to a much smaller extent. Including additional Rydberg states is particularly demanding for active-space-based methods and, as noted earlier, might be highly sensitive to the choice of basis set.
@@ 2.4 Calculation of observables
This comparative analysis allows us to stress some of the final findings of this exercise. First of all, most contributions employing an electronic-structure method including static correlation obtained in general a good description of most features of the , yet with significant differences in the evolution timescales for the main positive and negative features of this observable. XMS-CASPT2-based nonadiabatic molecular dynamics simulations lead, overall, to a better agreement with experiment for the overall trends and lifetime of features appearing on the map (e.g., predicted signals in Fig. 7 for Papers 2, 8, 12). Interestingly, some inconsistency between MCSCF methods can be observed: predictions 7, 10 and 15, do not exhibit the same features as the rest of the predictions employing MCSCF methods (for Paper 15, instabilities in the active space were noted). This observation further highlights that MCSCF methods, namely SA-CASSCF in the present case, are not black box tools and specific settings of parameters, such as active space or number of states, strongly determine the outcome of the calculations. The predictions based on single-reference methods (predicted signals in Fig. 7 for Papers 1, 3, 11) mostly exhibit a nearly static signal emanating from dynamics near the S2(3s) excited-state minimum, lacking the dissociation features observed with predictions using a method including static correlation. Despite the apparent discrepancies between the predicted signals represented as colormaps, the reader should bear in mind that the quantitative analysis offered by most contributions (and depicted in Fig. 8 with lineouts) is already useful to understand the combination of processes occurring during the photodynamics of cyclobutanone.
@@
While the maps plotted in Fig. 7 are useful to depict the overall time-evolution of the UED signal, they do not permit a clear evaluation of the temporal evolution of their respective features. To unravel the time-dependence of key features in the signal, we provide lineouts at and in Fig. 8, using both absolute (Figs. 8A and C) and normalized (Figs. 8B and D) intensities. All lineouts are integrated in a range of to reduce experimental noise and mimic the standard analysis of an experimental signal. The normalized data (normalized with respect to the maximum value along each lineout) are provided here only to facilitate a comparison with the experimental data, but the comparison between absolute scale and normalized data highlights the danger of such a comparison (see end of this paragraph). The lineout at is connected to the ring opening of cyclobutanone, while that at is connected to the fragmentation of the molecule and formation of photoproducts. Fig. 8B reinforces the observation of the previous paragraph that predictions employing an electronic-structure method including only static correlation predict too fast dynamics, whereas those making use of an approach combining static and dynamic correlation provide a closer match with the experimental time scale. This dramatic impact of the electronic-structure method on the outcome of the nonadiabatic molecular dynamics is clearly highlighted by Fig. 9A (and Fig. 9C), which represents the same lineouts as in Fig. 8 but now with each contribution colorcoded based on the type of correlation included in the electronic-structure method employed for the nonadiabatic dynamics (dynamic correlation, static correlation, dynamic+static correlation). Fig. 9A clearly highlights the trends (discussed earlier for the ring-opening process with a few exceptions): electronic-structure methods including solely dynamic correlation (yellow curves) lead to a slower adiabatic ring-opening process in S2 in comparison to nonadiabatic dynamics using approaches including both static and dynamic correlation (purple curves); methods including static correlation only (cyan curves), on the contrary, induce a much faster ring-opening dynamics than that observed with methods including both static and dynamic correlation. The signal obtained in Paper 14 appears as an outlier, which is possibly connected to the fact that MRCIS includes less dynamic correlation than the MRPT approaches. However, a correct timescale for a feature of the signal does not mean a correct prediction: Fig. 8A shows that the predictions can have significant variations in signal intensities, highlighting an important variation of the overall population of trajectories leading to a given signal. For example, the time trace of the normalized lineout at predicted by Paper 10 using MCSCF appears to have a reasonable timescale in Fig. 8B, but this feature is associated with a nearly zero intensity in the absolute signal (Fig. 8A). The initial increase of the experimental signal at (Fig. 8C and Fig. 8D) is due to the bond-elongation event during ring opening and dissociations. The following negative signal at later times is caused by the formation of photoproducts, which lack the signal associated with larger bond distances observed in cyclobutanone between carbons and oxygen. This feature is captured only by three predictions, namely Paper 6 (SA-CASSCF), 12 (XMS-CASPT2), and 14 (MRCIS), which simulated the ground-state dynamics of cyclobutanone and its photoproducts for a longer timescale. Finally, and keeping in mind the biased distribution towards mixed quantum/classical methods during this challenge, it is interesting to note that the trends observed in the lineouts do not seem to be correlated with the family of methods used for the nonadiabatic molecular dynamics (Fig. 9B and Fig. 9D). Additional simulations would be required to further support this lack of correlation.